An investment project provides cash inflows of $680 per year for 13 years.
a. What is the project payback period if the initial cost is $2,720?

What is the project payback period if the initial cost is $5,032?

What is the project payback period if the initial cost is $9,520?
Answer #1

a)

Project payback = Initial cost / cash flows

Project payback = 2,720 / 680

Project payback = 4 years

b)

Project payback = Initial cost / cash flows

Project payback = 5,032 / 680

Project payback = 7.4 years

c)

Project payback = Initial cost / cash flows

Project payback = 9,520 / 680

Project payback = 14 years